Part 1: Optimal order quantity

Joshua is running a pharmacy in Nigeria (where Naira is the local currency). His recent experience in the MicroMasters course motivated him to rethink how to reorder pharmaceuticals to replenish inventory. He recalls from Module 4, Unit 1, Video 5, that he should consider the purchase cost, the order cost, and the holding cost. In the past, Joshua rigorously tracked these cost items. For Ibuprofen he knows that:

unit cost: c = 368 Naira/pack

demand: D = 670 packs/year

ordering cost: A = 2300 Naira/order

holding cost: r∗c = 47 Naira/pack*year

Part 1 Answer: How many packs of Ibuprofen should Joshua order to minimize the total cost?

256.07 Correct

Part 2: Consolidating orders

Joshua runs a second pharmacy on the other side of town. He has been ordering for each location individually from a wholesaler. He realizes that he can save some money by consolidating the two orders into one single order. The costs per pack for both pharmacies are exactly the same (use A, c and r∗c values from Part 1). The annual demand for Ibuprofen in the second pharmacy is 1150 packs.

Part 2 Answer: What is the consolidated optimal order quantity (in packs of Ibuprofen)?

Part 3: Savings from consolidation

Part 3 Answer: If Joshua places one consolidated order (as in Part 2) instead of placing two separate orders, how much money will he save?
Answer #1

Part 1 Answer: = EOQ= (2*D*S/H)0.5 = (2*670*2300/47)0.5 =256.07

Part 2 Answer: What is the consolidated optimal order quantity (in packs of Ibuprofen)?

Answer= EOQ= (2*D*S/H)0.5 = (2*(670+1150)*2300/47)0.5 =422.05

Part 3 Answer: If Joshua places one consolidated order (as in Part 2) instead of placing two separate orders, how much money will he save?

EOQ when only 1150 units are placed = (2*1150*2300/47)0.5 =335.49

Total cost for placing 256.07 units = (256.07*47)/2 +(670/256.07)*2300 =12035.53

Total cost for placing 335.49 units = (335.49*47)/2 +(1150/335.49)*2300 =15768.01

Total cost for placing 422.05 units = (422.05*47)/2 +(1820/422.05)*2300 =19836.43

Saving =12035.53+15768.01-19836.43

Saving=7967.11 Naira